Peter has an annual salary of $29,523.60 for his second year on the job. If this is 6.2% more than his first year salary, how much did he earn his first year?
Answer by jmg(22)

This is kind of hard to explain unless you are good with percents.
He got a 6.2% raise. That means he makes what he used to make (100%) plus the 6.2%.
So now he is making 100% + 6.2% = 106.2% is what he is making now
You want to know 106.2% of what number is $29523.60.
If equation form,
106.2% * n = 29523.60 [change percent to decimal: 2 places left]
1.062 * n = 29523.60 so divide by 1.062 on both sides
n = $27,800 --> this is your answer


Double check. 27800 * 6.2% = 27800 * 0.062 = 1723.60 would be the raise.
Add that onto 27800 and you get 29523.60.